OpenStudy (anonymous):

Consider the absolute value inequality |x - 3| > 2. Part 1: Using complete sentences, explain whether this inequality will be an "and" compound inequality (conjunction) or an "or" compound inequality (disjunction) and why. Part 2: Provide the solution to the inequality and describe the graph in complete sentences.

OpenStudy (kirbykirby):

|x| = -x if x < 0 = x if x \(\ge\) 0 You can extend this to |x-3| = -(x-3) if x-3 < 0 = x-3 if x-3 \(\ge\) 0

OpenStudy (kirbykirby):

and take this further using the inequality So you have to solve x-3 > 2 and -(x-3) > 2

OpenStudy (anonymous):

what? you lost me.... :(

OpenStudy (kirbykirby):

ok if you had |x-3| = 2, do you know how to solve that?

OpenStudy (kirbykirby):

because then solving |x-3| > 2 is almost the same thing, but with minor details added

OpenStudy (anonymous):

yea you add the three to the other side

OpenStudy (anonymous):

hold on ill be back on in 15 mins

OpenStudy (kirbykirby):

ok

OpenStudy (kirbykirby):

What you wrote is partially correct. When you solve an absolute value, you are really solving 2 equations